OpenStudy (anonymous):

Solve for x. 9 = x/2 + 1

pooja195 (pooja195):

subtract 1 from both sides 9-1=?

OpenStudy (anonymous):

i like you bc you teach me d=not tell me answer

OpenStudy (anonymous):

8

pooja195 (pooja195):

thanks :) and goo 8=x/2 we multiply by the reciprocal so x/2=2/x 8 x 2 =?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Oops wrong key

pooja195 (pooja195):

good so x=16

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yep x=16